What is the method of obtaining the moment of inertia of a cylinder of mass M and radius R and length L through


(i) the centre perpendicular to the length


(ii) end faces and perpendicular to the length.


Please explain.

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Hinan ali

we know that

  Iz= mr2/2

and Ix + Iy =Iz

by symmetry Ix =Iy

so Ix =Iz/2

        =mr2/4


and from parallel axix theorem  Ix' = Ix + mx2

  so                                         Ix' = Ix + mL2/4

                                                    =mr2/4 + mL2/4


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin